# FedCQA: Answering Complex Queries on Multi-Source Knowledge Graphs via Federated Learning

### v2

@@ 3.2. Complex Logical Query
The complex knowledge graph query is defined in existential positive first-order logic form, consisting of various types of logic expressions like existential quantifiers $\exists$, logic conjunctions $\wedge$, and disjunctions $\vee$. In the logical expression, there is a set of anchor entities $V_{a}\in\mathcal{V}$ denotes given context, existential quantified variables $V_{1},V_{2},...,V_{k}\in\mathcal{V}$, and a unique variable $V_{?}$ denotes our query target. The complex query intends to find the target answers $V_{?}\in\mathcal{V}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the knowledge graph that can satisfy the given logical expressions simultaneously. Following the definition in (Ren et al., 2020), the complex query expression can be converted to the disjunctive normal form (DNF) in the following:
@@
Figure 2. The training and retrieval process of FedCQA. The blue line denotes the training process, and the green line denotes the retrieval process. In the training, clients’ CQA models are trained on respective knowledge graphs. At each round, local CQAs are aggregated at the server and updated using the global parameters. Among them, the embeddings are protected by secret aggregation so that the server can not access them. In the retrieval, each query is decomposed into sub-queries. Clients compute sub-query embeddings which the server are used to aggregate query embeddings. Answer scores are computed at clients and are aggregated at server to retrieve answers.
3.3. Knowledge Graph Set Query
Knowledge graphs are owned by different data holders and cannot be shared directly with each other, therefore, a complex query $q$ can involve entities and relations from different knowledge graphs. We define those queries as follows:
@@ 3.3. Knowledge Graph Set Query
A complex query $q$ is a cross-graph query if there exists query answers $V_{?}\in\mathcal{V}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the knowledge graph that can satisfy the given logical expressions and the atomic expressions in the query can not be found in a single knowledge graph.
For example in figure 1, the entity ”Physics” is the answer to the query $q$ because there exists an existentially quantified variable ”Einstein” that can satisfy the logical expression. The query $q$ is a cross-graph query as the atomic expressions in the query are from different knowledge graphs and can not be found in a single graph. For example, Win(Einstein, Novel Prize) and BornIn(Einstein, Germany) are two atomic expressions from different knowledge graphs. When the distributed knowledge graph set faces cross-graph queries, the answer can not be inferred from one single knowledge graph. Besides, We have the definition of in-graph query correspondingly:
@@
A complex query $q$ is an in-graph query if for all answers $V_{?}\in\mathcal{V}$ to the query, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the knowledge graph that can satisfy the given logical expressions and the atomic expressions in the query are from a single knowledge graph.
These in-graph queries can retrieve answers according to one of the knowledge graphs in the KG set and can be solved with existing complex query answering models. However, in a distributed knowledge graph complex query answering system, the in-graph queries may retrieve more answers as more knowledge are provided.
@@ 3.4. Problem Formulation
Given a knowledge graph set $\mathcal{G}=\{g_{1},...,g_{N}\}$ with $N$ knowledge graphs. Every knowledge graph is owned by independent data holders and can not be shared to construct a unified knowledge graph, which means that the triples in each knowledge graph will stay private at local devices. The knowledge graphs in the $\mathcal{G}$ are related and have part of entities and relation overlapped. There are complex queries involving elements from the knowledge graph set $\mathcal{G}$ and can be classified as cross-graph queries and in-graph queries. We aim to construct a distributed complex query answering system to reason over multi-source knowledge graphs and retrieve answers to the queries while keeping privacy preserved. To achieve this, we assume that there is an honest but curious server managing the federated complex query answering system. Because the learned embeddings are vulnerable to various privacy attacks, the embeddings on the server should be further protected.
4. Federated Complex Query Answering
In this section, we introduce the learning and query retrieval process of our proposed FedCQA.
@@ 4.1. Model Learning
We first introduce the training of FedCQA. As shown in Figure 2, FedCQA has a central server and a set of clients. Each client has a knowledge graph with overlapping entities of others. The server takes the responsibility of aggregating parameters and organizing the training and retrieval process. The clients train a local CQA model based on their knowledge graph. According to the sensitivity of the parameters, we divide the query encoding methods into two parts: operator function with parameter $\Theta$ and entity embeddings $\mathbf{E}$. For the operator function, the client directly sends the parameter to the server for aggregation and receives the global function to update the local operator function, which is similar to FedAvg (McMahan et al., 2017). Therefore, in the following parts, we only introduce the entity embeddings aggregation in FedCQA.
@@ 4.1.1. Secret Aggregation
There are various techniques, like homomorphic encryption (HE) (Paillier, 1999), secure multi-party computation (MPC) (Mohassel and Zhang, 2017), and differential privacy (Dwork, 2008; Geyer et al., 2017) to protect the uploaded parameters, however the protection of aggregated global model are often ignored. Unfortunately, the global model is informative and vulnerable to privacy attacks  (Zhang et al., 2022). Hence we propose a secret aggregation applied in parameter aggregation which can prevent the global server from knowing the aggregated parameters. Assume that at each client $i$, a parameter denoted as $\theta_{i}$ is uploaded to a server for aggregating global parameter $\theta$. The procedure of secret aggregation is described in Algorithm 1. In the beginning, each client $C_{i}$ randomly generates parameters $\theta_{i}^{r}$ and shares it with other clients with encryption (Shown in Appendix A). Each client has a set of parameters $\{\theta_{1}^{r},\theta_{2}^{r}.\cdots,\theta_{n}^{r}\}$. For each client, $C_{i}$ uploads perturbed parameter $(\theta_{i}+\theta_{i}^{r})$ to server. The server collects all perturbed parameters from clients, computes aggregated perturbed parameter $\theta^{r}$, and sends it back to all clients. Finally, Clients can compute aggregated parameter $\theta$ and prevent from exposing it to the server.
0:  $n$ clients $C_{1},C_{2},\ldots,C_{n}$, client $C_{i}$ has parameter $\theta_{i}$
  Each client $C_{i}$ generates random parameter $\theta_{i}^{r}$
1  Transmit $\theta_{i}^{r}$ to all other clients with encryption, each
client has a set of parameters $\{\theta_{1}^{r},\theta_{2}^{r}.\cdots,\theta_{n}^{r}\}$
  Client $C_{i}$:
  Upload perturbed parameter $(\theta_{i}+\theta_{i}^{r})$ to server
  Receive parameters $\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})$ from server
  Compute parameters $\theta=[\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})-\sum_{i}^{n}\theta_{i}^{r}]/n$
  Server:
  for $i=1,\cdots,n$ do
     Receive parameters $(\theta_{i}+\theta_{i}^{r})$ from client $C_{i}$
  end for
  Compute parameters $\theta^{r}=\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})$
  Send parameters $\theta^{r}$ to all clients
Algorithm 1 Secret Aggregation
@@ 4.1.2. Model Training
FedCQA performs secret aggregation for entity embeddings. First, the client $C_{i}$ will randomly initialize the local entity embeddings $\mathbf{E}_{0}^{i}\in\mathbb{R}^{n_{i}\times d}$ and perturbation embeddings $\mathbf{E}_{r}^{i}\in\mathbb{R}^{n_{i}\times d}$. Every client will share the permutation embeddings with all clients. At round $t$, the server will select part of the clients $\mathbf{C}$ participating in the training. After local training of CQA on respective local knowledge graphs, client $C_{i}$ sends perturbed local entity embeddings $\mathbf{E}_{t}^{i}+\mathbf{E}_{r}^{i}$ to the server. The server will aggregate the entity embeddings using perturbed local embeddings:
@@
After secret aggregation, the entity embeddings of all clients are shared without exposing the sensitive information to the server. Besides the training of entity embeddings, the operator networks are trained and aggregated using FedAvg (McMahan et al., 2017), and the detailed descriptions of the FedCQA are shown in Appendix B
@@ 4.2. Query Retrieval
After training, the server in FedCQA manages the process of retrieving answers to complex queries.
@@ 4.2.1. Query Encoding
Query encoding methods commonly represent queries to embeddings and retrieve answers according to scoring functions where similarity functions are widely used. FedCQA encodes queries and treats two types of queries differently. For in-graph queries, FedCQA directly encodes queries using corresponding local complex query answering models. For cross-graph queries, as the query involves entities from multiple knowledge graphs, we decompose the queries to atomic expressions and encode these atomic queries using local CQA models, and use global operator function models to compute the representations of the original queries.
@@ 4.2.2. Answer Retrieval
Because the entity embeddings are not stored in the central server, we can only retrieve answers from all distributed local knowledge graphs. Given a query encoding $q$, we score all the candidate entities at each local knowledge graph, at client $C_{i}$:
@@
where $f^{i}_{s}$ is a score function in the client $C_{i}$. Then the score will be uploaded to the server to aggregate a score table for all unique entities in the knowledge graph sets:
@@
The final answers to the queries are retrieved globally from the knowledge graph set according to the score table.
@@ 5. Experiments
In this section, we create a benchmark for distributed knowledge graph complex query answering problems and evaluate our proposed FedCQA’s performance on the benchmark.
@@ 5.1.1. Datasets
In our experiment, following previous work, we use the three commonly used knowledge graph datasets: FB15k (Bollacker et al., 2008; Bordes et al., 2013), FB15k-237 (Toutanova and Chen, 2015), and NELL995 (Carlson et al., 2010) to construct the distributed query answering benchmark. In each dataset, there are vertices describing entities and edges describing relations. To evaluate the distributed complex query, we conduct experiments assuming having 3 and 5 clients in each federated complex query answering system, respectively. We randomly select relations into clients and split triples into clients according to selected relations as a local knowledge graph. The triplets in each local knowledge graph are separated into training, validation, and testing with a ratio of 8:1:1 respectively. Following previous works (Ren et al., 2020), we construct training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{val}$, and test graph $\mathcal{G}_{test}$ in each client by training edges, training+validation edges, and training+validation+testing edges, respectively. The detailed statistics are listed in the table 1. #Clients denotes the number of clients, #Nodes, #Relations, #Edges denote the average number of nodes, relations, and edges in each client respectively.
@@ 5.1.2. Query Sampling
Following previous work (Hamilton et al., 2018; Bai et al., 2023b), we evaluate the complex query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,ip,3i,pi,2u$, and $up$. As shown in figure 3, each subgraph denotes one query type, where each edge represents either a projection or a logical operator, and each node represents either a set of entities, the anchor entities and relations are to be specified to instantiate logical queries. We use the sampling method commonly used in previous works (Bai et al., 2023c; Ren et al., 2020) to randomly sample complex queries from knowledge graphs. We randomly sample two sets of queries from the knowledge graph sets: in-graph queries and cross-graph queries to evaluate local and global answer retrieval performance. For local model evaluation, we first obtain training, validation, and testing queries from the formerly constructed local knowledge graphs respectively. Then for the training queries, we conduct a graph search to find corresponding training answers on the local training graph. For the validation queries, we search for the answers on both the training graph and the validation graph and only use those queries that have different numbers of answers on the two graphs. For the testing queries, we use those queries that have different answers on the testing graph from answers on the training graph and validation graph. For global model evaluation, we construct global training, validation, and testing knowledge graphs using all local knowledge graphs, and sample testing queries with atomic expressions from different local knowledge graphs, finally we search for answers on three global knowledge graphs and only use those queries that have different answers on the testing graph from other two graphs. We collect statistics of complex queries in three datasets and the statistics are shown in Table 2. The number of in-graph queries is the average number of the client’s local queries.
@@
Table 3. The retrieval performance of distributed knowledge graph complex query answering models when there are 3 clients. The average results of HR@3 and MRR of all clients are reported. The best results are underlined. The best results of distributed models are in bold.
@@ 5.1.3. Baselines
We can use various existing query encoding methods as our local base model, to evaluate the effectiveness and generalization ability of our proposed FedCQA, we select three commonly used complex encoding methods GQE (Hamilton et al., 2018), Q2P (Bai et al., 2022), tree-LSTM (Bai et al., 2023c) as our base model. GQE is a graph query encoding model that encodes a complex query into a vector in embedding space; Q2P represents complex queries using multiple vectors; tree-LSTM recursively represents complex queries and treats all operations, entities, and relations as tokens.
To the best of our knowledge, there are no existing federated complex query answering methods but several federated knowledge graph embedding methods, therefore, we choose to compare our methods with FedE (Chen et al., 2021) and FedR (Zhang et al., 2022), two commonly used federated knowledge graph embedding methods as baselines. FedE aggregates both entity embeddings and relation embeddings in a server, while FedR only aggregates relation embeddings for privacy concerns and communication efficiency. We utilize these two methods with slight modifications to train a global complex query answering model: FedE aggregates all query encoding parameters and FedR aggregates relation embeddings and query encoding networks. Besides, we also compare our FedCQA with local and central settings. In the local setting, there is no collaboration between clients, while in the central setting, all distributed graphs in the knowledge graph set are aggregated for a global graph for training, we sample complex queries from global training, and validation knowledge graphs for training and validation.
If there is no further statement, we use the following implementation settings in the experiments. We evaluate FedCQA on three base models: GQE, Q2P, and tree-LSTM in FedE, FedR, and our proposed FedCQA for comparison. We tune hyper-parameters on the validation local queries for the base query encoding methods and set the dimension of entities and relations as 400 for all models for fair comparison and use AdamW (Loshchilov and Hutter, 2018) as optimizer.
Table 4. The retrieval performance of distributed knowledge graph complex query answering models when there are 5 clients. The average results of HR@3 and MRR of all clients are reported. The best results are in bold.
FB15k-237 FB15k NELL
@@
tree-LSTM FedCQA 13.79 13.27 12.74 12.18 15.44 15.81 15.28 14.24 15.68 14.28 14.57 12.89

### v3

@@ 3.2. Complex Logical Query
The complex logical query is defined in existential positive first-order logic form, consisting of various types of logic expressions like existential quantifiers $\exists$, logic conjunctions $\wedge$, and disjunctions $\vee$. In the logical expression, there is a set of anchor entities $V_{a}\in\mathcal{V}$ denotes given context, existential quantified variables $V_{1},V_{2},...,V_{k}\in\mathcal{V}$, and a unique variable $V_{?}$ denotes our query target. The complex query intends to find the target answers $V_{?}\in\mathcal{V}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the graph-structured data that can satisfy the given logical expressions simultaneously. Following the definition in (Ren et al., 2020), the complex query expression can be converted to the disjunctive normal form (DNF) in the following:
@@
Figure 2. The training and retrieval process of FedNGDB. The blue line denotes the training process, and the green line denotes the retrieval process. In the training, clients’ NGDB models are trained on respective graph-structured data. At each round, local NGDBs are aggregated at the server and updated using the global parameters. Among them, the embeddings are protected by secret aggregation so that the server can not access them. In the retrieval, each query is decomposed into sub-queries. Clients compute sub-query embeddings which the server is used to aggregate query embeddings. Answer scores are computed at clients and are aggregated at a server to retrieve answers.
3.3. Distributed Graph Set Query
Graph databases are owned by different data holders and cannot be shared directly with each other, therefore, a complex query $q$ can involve entities and relations from different graphs. We define those queries as follows:
@@ 3.3. Distributed Graph Set Query
A complex query $q$ is a cross-graph query if there exists query answers $V_{?}\in\mathcal{V}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the graph that can satisfy the given logical expressions and the atomic expressions in the query can not be found in a single graph.
For example in figure 1, the entity ”Physics” is the answer to the query $q$ because there exists an existentially quantified variable ”Einstein” that can satisfy the logical expression. The query $q$ is a cross-graph query as the atomic expressions in the query are from different graph databases and can not be found in a single graph. For example, Win(Einstein, Nobel Prize) and BornIn(Einstein, Germany) are two atomic expressions from different graph databases. When the distributed graph databases face cross-graph queries, the answer can not be inferred from a single graph. Besides, We have the definition of in-graph query correspondingly:
@@
A complex query $q$ is an in-graph query if for all answers $V_{?}\in\mathcal{V}$ to the query, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the graph that can satisfy the given logical expressions and the atomic expressions in the query are from a single graph database.
These in-graph queries can retrieve answers according to one of the graphs in the graph set and can be solved with existing complex query answering models. However, in a distributed graph neural graph database system, the in-graph queries may retrieve more answers as more knowledge is provided.
@@ 3.4. Problem Formulation
Given a graph-structured data set $\mathcal{G}=\{g_{1},...,g_{N}\}$ with $N$ graphs. Every graph is owned by independent data holders and can not be shared to construct a unified graph database. We assume that the triples are sensitive as they describe the informative relations between entities while the index of entities and relations can be shared, which means that the triples in each graph database will stay private on local devices. The graphs in the $\mathcal{G}$ are related and have part of entities and relation overlapped. There are complex queries involving elements from the graph set $\mathcal{G}$ and can be classified as cross-graph queries and in-graph queries. We aim to construct a distributed neural graph database system to reason over multi-source graphs and retrieve answers to complex logical queries while keeping privacy preserved, especially the triple information indicating the relations between entities. To achieve this, we assume that there is an honest but curious server managing the federated neural graph database system. Because the learned embeddings are vulnerable to various privacy attacks, the embeddings cannot be exposed to the server and should be further protected before being transferred to the server.
4. Federated Neural Graph Databases
In this section, we introduce the learning and query retrieval process of our proposed FedNGDB.
@@ 4.1. Model Learning
We first introduce the training of FedNGDB. As shown in Figure 2, FedNGDB has a central server and a set of clients. Each client has a graph with overlapping entities of others. The server takes the responsibility of aggregating parameters and organizing the training and retrieval process. The clients train a local NGDB model based on their graph-structured data. According to the sensitivity of the parameters, we divide the query encoding methods into two parts: operator function with parameter $\Theta$ and entity embeddings $\mathbf{E}$. For the operator function, the client directly sends the parameter to the server for aggregation and receives the global function to update the local operator function, which is similar to FedAvg (McMahan et al., 2017). Therefore, in the following parts, we only introduce the entity embeddings aggregation in FedNGDB.
@@ 4.1.1. Secret Aggregation
There are various techniques, like homomorphic encryption (HE) (Paillier, 1999), secure multi-party computation (MPC) (Mohassel and Zhang, 2017), and differential privacy (Dwork, 2008; Geyer et al., 2017) to protect the uploaded parameters, however the protection of aggregated global model are often ignored. Unfortunately, the global model is informative and vulnerable to privacy attacks  (Zhang et al., 2022; Hu and Song, 2023). Hence we propose a secret aggregation applied in parameter aggregation which can prevent the global server from knowing the aggregated parameters using homomorphic encryption. Assume that at each client $i$, a parameter denoted as $\theta_{i}$ is uploaded to a server for aggregating global parameter $\theta$. The procedure of secret aggregation is described in Algorithm 1. In the beginning, each client $C_{i}$ randomly generates perturbed parameters $\theta_{i}^{r}$ and shares them with other clients with encryption (Shown in Appendix A). After the sharing, each client has a set of parameters $\{\theta_{1}^{r},\theta_{2}^{r}.\cdots,\theta_{n}^{r}\}$. In the training process, for each client, $C_{i}$ uploads perturbed parameter $(\theta_{i}+\theta_{i}^{r})$ to the server under the protection of homomorphic encryption. The server collects all perturbed parameters from clients, computes aggregated perturbed parameter $\theta^{r}$, and sends it back to all clients. Finally, after receiving the perturbed parameters, the clients first decrypt the parameters and can compute aggregated parameter $\theta$ by removing the perturbed parameters and preventing exposing it to the server.
@@ 4.1.2. Model Training
FedNGDB performs secret aggregation for entity embeddings. First, the client $C_{i}$ will randomly initialize the local entity embeddings $\mathbf{E}_{0}^{i}\in\mathbb{R}^{n_{i}\times d}$ and perturbation embeddings $\mathbf{E}_{r}^{i}\in\mathbb{R}^{n_{i}\times d}$. Every client will share the permutation embeddings with all clients. At round $t$, the server will select part of the clients $\mathbf{C}$ participating in the training. After local training of CQA on respective local graphs, client $C_{i}$ sends perturbed local entity embeddings $\mathbf{E}_{t}^{i}+\mathbf{E}_{r}^{i}$ to the server. The server will aggregate the entity embeddings using perturbed local embeddings:
@@
After secret aggregation, the entity embeddings of all clients are shared without exposing the sensitive information to the server. Besides the training of entity embeddings, the operator networks are trained and aggregated using FedAvg (McMahan et al., 2017), and the detailed descriptions of the FedNGDB are shown in Algorithm 2.
@@ 4.2. Query Retrieval
After training, the server in FedNGDB manages the process of retrieving answers to complex queries. The server first tries to arrange related clients to encode the coming queries and retrieves answers from all local graph databases based on the encoding.
@@ 4.2.1. Query Encoding
Query encoding methods commonly represent queries to embeddings and retrieve answers according to scoring functions where similarity functions are widely used. FedNGDB encodes queries and treats two types of queries differently. For in-graph queries, as these queries only involve a single graph data, FedNGDB can directly encode queries using corresponding local complex query answering models. For cross-graph queries, as the query involves entities from multiple graphs, the server will take the responsibility to plan the entire encoding process: First, the server will decompose the query to atomic expressions and send each expression to corresponding local graphs. The clients encode received atomic queries using their own local CQA models and send back the results to the server. The server collects all the encoding results and uses global operator function models to compute the representations of the queries. The query is iteratively updated by communicating between the server and clients until the original queries are encoded.
@@ 4.2.2. Answer Retrieval
Because the entity embeddings are not stored in the central server, we can only retrieve answers from all distributed local graphs after encoding the queries. Given a query encoding $q$, we score all the candidate entities at each local graph database, at client $C_{i}$:
@@
where $f^{i}_{s}$ is a score function in the client $C_{i}$. Then the score will be uploaded to the server to aggregate a score table for all unique entities in the graph sets:
@@
The final answers to the queries are retrieved globally from the graph database set according to the score table.
@@ 5. Experiments
In this section, we create a benchmark of distributed graph complex logical query answering problems for distributed neural graph databases and evaluate our proposed FedNGDB’s performance on the benchmark.
@@ 5.1.1. Datasets
In our experiment, following previous work, we use the three commonly used knowledge graphs as graph-structured data: FB15k (Bollacker et al., 2008; Bordes et al., 2013), FB15k-237 (Toutanova and Chen, 2015), and NELL995 (Carlson et al., 2010) to construct the distributed query answering benchmark. In each dataset, there are vertices describing entities and edges describing relations. To evaluate the distributed complex query, we conduct experiments assuming having 3 and 5 clients in each federated neural graph database system, respectively. We randomly select relations into clients and split triples into clients according to selected relations as a local graph database. The triplets in each local graph are separated into training, validation, and testing with a ratio of 8:1:1 respectively. Following previous works (Ren et al., 2020), we construct training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{val}$, and test graph $\mathcal{G}_{test}$ in each client by training edges, training+validation edges, and training+validation+testing edges, respectively. The detailed statistics are listed in the table 1. #Clients denotes the number of clients, #Nodes, #Relations, #Edges denote the average number of nodes, relations, and edges in each client respectively.
@@ 5.1.2. Query Sampling
Following previous work (Hamilton et al., 2018; Bai et al., 2023b), we evaluate the complex logical query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,ip,$ $3i,pi,2u$, and $up$. As shown in figure 3, each subgraph denotes one query type, where each edge represents either a projection or a logical operator, and each node represents either a set of entities, the anchor entities and relations are to be specified to instantiate logical queries. We use the sampling method commonly used in previous works (Bai et al., 2023c; Ren et al., 2020) to randomly sample complex queries from graphs. We randomly sample two sets of queries from the graph sets: in-graph queries and cross-graph queries to evaluate local and global answer retrieval performance. For local model evaluation, we first obtain training, validation, and testing queries from the formerly constructed local graph databases respectively. Then for the training queries, we conduct a graph search to find corresponding training answers on the local training graph. For the validation queries, we search for the answers on both the training graph and the validation graph and only use those queries that have different numbers of answers on the two graphs. For the testing queries, we use those queries that have different answers on the testing graph from answers on the training graph and validation graph. For global model evaluation, we construct global training, validation, and testing graphs using all local graphs, and sample testing queries with atomic expressions from different local graphs, finally we search for answers on three global graphs and only use those queries that have different answers on the testing graph from other two graphs. We collect statistics of complex queries in three datasets and the statistics are shown in Table 2. The number of in-graph queries is the average number of the client’s local queries.
@@
Table 3. The retrieval performance of distributed neural graph databases when there are 3 clients. The average results of HR@3 and MRR of all clients are reported. The best results are underlined. The best results of distributed models are in bold.
@@ 5.1.3. Baselines
We can use various existing query encoding methods as our local base model, to evaluate the effectiveness and generalization ability of our proposed FedNGDB, we select three commonly used complex encoding methods GQE (Hamilton et al., 2018), Q2P (Bai et al., 2022), Tree-LSTM (Bai et al., 2023c) as our base model. GQE is a graph query encoding model that encodes a complex query into a vector in embedding space; Q2P represents complex queries using multiple vectors; Tree-LSTM recursively represents complex queries and treats all operations, entities, and relations as tokens.
To the best of our knowledge, there are no existing federated complex query answering methods but several federated knowledge graph embedding methods, therefore, we choose to compare our methods with FedE (Chen et al., 2021) and FedR (Zhang et al., 2022), two commonly used federated knowledge graph embedding methods as baselines. FedE aggregates both entity embeddings and relation embeddings in a server, while FedR only aggregates relation embeddings for privacy concerns and communication efficiency. We utilize these two methods with slight modifications to train a global complex query answering model: FedE aggregates all query encoding parameters and FedR aggregates relation embeddings and query encoding networks. Besides, we also compare our FedNGDB with local and central settings. In the local setting, there is no collaboration between clients, while in the central setting, all distributed graphs in the graph set are aggregated for a global graph for training, we sample complex queries from global training, and validation graphs for training and validation.
If there is no further statement, we use the following implementation settings in the experiments. We tune hyper-parameters on the validation local queries for the base query encoding methods and set the dimension of entities and relations as 400 for all models for fair comparison and use AdamW (Loshchilov and Hutter, 2018) as optimizer.
Table 4. The retrieval performance of distributed neural graph databases when there are 5 clients. The average results of HR@3 and MRR of all clients are reported. The best results are in bold.
FB15k-237 FB15k NELL995
@@
Tree-LSTM FedNGDB 13.79 13.27 12.74 12.18 15.44 15.81 15.28 14.24 15.68 14.28 14.57 12.89
5.1.4. Evaluation Metrics
Following the previous work (Bai et al., 2023b), we evaluate the generalization capability of models by calculating the rankings of answers that cannot be directly retrieved from an observed graph. Given a testing query $q$, the training, validation, and public testing answers are denoted as $M_{train}$, $M_{val}$, and $M_{test}$, respectively. We evaluate the quality of retrieved answers using Hit ratio (HR) and Mean reciprocal rank (MRR). HR@K metric evaluates the accuracy of retrieval by measuring the percentage of correct hits among the top K retrieved items. The MRR metric evaluates the performance of a ranking model by computing the average reciprocal rank of the first relevant item in a ranked list of results. The metric can be defined as:
$m(r)=\textbf{1}[r\leq K]$ if the metric is HR@K and $m(r)=\frac{1}{r}$ if the metric is MRR. Higher values denote better reasoning performance. We train local models at each client by using the in-graph training queries and tune hyper-parameters using the validation queries. The evaluation is then finally conducted on the testing queries, including the evaluation of in-graph queries on local query encoding models and cross-graph queries on the global federated neural graph database system, respectively.